Cost Definitions and Calculations

For each of the following independent situations, calculate the missing values:

1. The Belen plant purchased $80,100 of direct materials during June. Beginning direct materials inventory was $2,700, and direct materials used in production were $73,900. What is ending direct materials inventory?
$

2. Forster Company produced 17,900 units at an average cost of $5.90 each. The beginning inventory of finished goods was $4,484. (The average unit cost was $5.90.) Forster sold 18,020 units. How many units remain in ending finished goods inventory?
units

3. Beginning work in process (WIP) was $115,200, and ending WIP was $119,800. If total manufacturing costs were $332,500, what was the cost of goods manufactured?
$

4. If the conversion cost is $92 per unit, the prime cost is $53, and the manufacturing cost per unit is $114, what is the direct materials cost per unit?
$ per unit

5. Total manufacturing costs for August were $411,000. Prime cost was $52,000, and beginning WIP was $80,000. The cost of goods manufactured was $448,000. Calculate the cost of overhead for August and the cost of ending WIP.

Cost of overhead for August $
Ending WIP $
Answer #1

Solution 1:

ending direct materials inventory =  Purchase of direct materials + Beginning direct materials inventory -  direct materials used in production

= $80100 + $2700 - $73900 = $8,900

Solution 2:

Units in ending finished goods inventory = produced units + beginning inventory Units - sold units.

= 17900 + ($4484/$5.90) - 18020

= 17900 + 760 - 18020

= 640 units

Solution 3:

cost of goods manufactured = Beginning WIP + Manufacturing costs - Ending WIP

= $115200 + 332500 - 119800 = $327,900

Solution 4:

Direct materials cost per unit =  manufacturing cost per unit - Conversion cost per unit

= $114 - $92 = $22 per unit

Solution 5:

Cost of overhead for August = Total Manufacturing costs - Prime costs

= 411000 - 52000 = $359,000

Cost of ending WIP = Total Manufacturing costs + Beginning WIP - cost of goods manufactured

= 411000 + 80000 - 448000 = $43,000